OpenStudy (anonymous):

Which expression is equivalent to Which expression is equivalent to 5(c + b) + 4c - 2b? a.)3c + 9b b.)5c + 3b c.)9c + 3b d.)9c + 7c

OpenStudy (anonymous):

5(c + b) + 4c - 2b = 5c+ 5b +4c-2b = 5c+ 4c +5b-2b= 9c+3b Hence option c.)9c + 3b is the correct answer..

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@adventuretimefreak7 you are most welcome...